OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8> 3.0° 6.0° -7.0° 16.0°

OpenStudy (anonymous):

You can use the dot product formula

OpenStudy (anonymous):

Or the cross product formula as well if you like

OpenStudy (anonymous):

\[ \cos\theta = \frac{u \cdot v}{\|u\|\|v\|} \]

OpenStudy (anonymous):

@juanjohnguy Show me how to cross product these vectors.

OpenStudy (anonymous):

hm...just realized that its only in R 2...not sure if it works in R 2

OpenStudy (anonymous):

But I was assuming we could just use the formula magnitude(u cross v) = mag(u) times mag(v) sin(theta)

OpenStudy (anonymous):

ok?